OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 124 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 28, how many women auditioned? 40 44 70 80

OpenStudy (anonymous):

First set up a system of equations to solve the problem: m + w = 124 2m - 3w = 28 Solve via elimination: Multiply the top equations by -2 -2m - 2w = -248 2m - 3w = 28 Add the equations together: -5w = -220 Divide both sides by -5. w = 44 Substitute 44 for w in the first equation: m + 44 = 128 Subtract 44 from both sides: m = 84 So 44 women auditioned and 84 men auditioned.

OpenStudy (anonymous):

Thank you! That helped a lot!